what is the diagonal measurement of a rectangle 12' x 16'
Answer by checkley77(12844)

The diagonal forms a right triangle of a 12,16,x triangle.
12^2+16^2=c^2
144+256=c^2
400=c^2
c=sqrt400
c=20 ans for the hypotenuse.